OpenStudy (anonymous):

Solve for a. a+1/4=a+2/6+a+3/9=2 0 7 9 63

OpenStudy (espex):

I would suggest pushing all the 'a' terms to the right and all the fractions to the left.

OpenStudy (espex):

Although I must say that the second = sign does puzzle me a bit.

OpenStudy (anonymous):

7

OpenStudy (anonymous):

?

OpenStudy (espex):

How does the second equal sign fit into the equation though, because a+1/4 is most likely not equal to 2.

OpenStudy (anonymous):

7+1 divided by 4 is 2

OpenStudy (espex):

Except that it would be 29/4 as it is not (7+1)/4

OpenStudy (espex):

The way it is written they all equal 2, and you cannot add 1/4 to a and get 2, at the same time add 2/6, 3.9, AND 'a' and get 2. Could you verify that the original equation is accurate.

OpenStudy (anonymous):

sorry i messed up its suppose to be 62 CoreyContini Solve for a. a+1/4-a+2/6+a+3/9=2

OpenStudy (espex):

AH.. that makes a big difference. :)

OpenStudy (maheshmeghwal9):

ans is 7.as u can take variables one side &constant one side

OpenStudy (espex):

Under the circumstances I'm not sure how you get any of the answers they list:\[a+\frac{1}{4}-a+\frac{2}{6}+a+\frac{3}{9}=2\rightarrow a+\frac{1}{4}+\frac{2}{6}+\frac{3}{9}=2\]\[a+\frac{1}{4}+\frac{2}{6}+\frac{3}{9}=2 \rightarrow a+\frac{9}{36}+\frac{12}{36}+\frac{12}{36}=2\]\[ a+\frac{9}{36}+\frac{12}{36}+\frac{12}{36}=2 \rightarrow a+\frac{33}{36}=2 \rightarrow a=2-\frac{33}{36}\]\[ a=2-\frac{33}{36} \rightarrow a=\frac{72}{36}-\frac{33}{36} \rightarrow a=\frac{39}{36} \rightarrow a=\frac{13}{12}\]